# Reassurance Robots: OCD in the Age of Generative AI

### v1

@@ Abstract.
Obsessive Compulsive Disorder (OCD) is a mental health disorder characterized by distressing repetitive patterns of thought, referred to as obsessions, and behaviors aimed to reduce the distress, referred to as compulsions. The explosion of artificial intelligence (AI) into the modern zeitgeist through the introduction of generative AI (GenAI) systems such as ChatGPT has led to novel obsessions and compulsions involving AI in individuals with OCD. Through a qualitative analysis of 100 Reddit posts related to AI on a popular subreddit for OCD, I explore the ways AI is impacting the presentation of OCD, including novel examples of AI-based obsessions and compulsions. I argue that GenAI in its current form harms individuals with OCD by becoming “Reassurance Robots,” and that future designs of GenAI must take OCD into account.
@@ 1. Introduction
Despite having a lifetime prevalence of 2.3% of the population in the United States (U.S.), with approximately 50% of those with the disorder reporting serious impairment (17), there is a major gap in the diagnosis and treatment of individuals with OCD. The mean difference between age of symptom onset and age of diagnosis is often over ten years (Ziegler et al., 2021), and a recent report found that up to 75% of cases of OCD in the United States are undetected or undiagnosed (Deusser et al., 2025), leaving a significant population with OCD symptoms that are currently untreated. As such, it is imperative that the disorder receive more attention and research to improve outcomes in individuals living with OCD.
With the advent of GenAI into the popular zeitgeist, it has been reported that GenAI has become a topic of obsessions and a site of compulsions for people with OCD (i.e. (Samuel, 2025)). Given the power to create and design technologies, the human-computer interaction (HCI) community must understand the harms that can be caused by GenAI to those with OCD. At a time when many researchers are looking to GenAI to help diagnose or treat OCD (Kim et al., 2025), there needs to be an understanding of how it can also cause harm. At this time, most research on the impact of GenAI on mental health has been in the context of depression and suicidality (Amine et al., 2025), however, given the high prevalence of OCD, it is critical that research study the impact of GenAI on OCD as well.
This paper contributes what I believe to be the first research exploring the ways in which GenAI has led to novel presentations of OCD through AI-based obsessions and compulsions. In the discussion, I argue that the current design of GenAI harms individuals with OCD by becoming “Reassurance Robots” and that future designs of GenAI should design defensively against this.
@@ 2. Background
OCD is often thought of as broken down into subtypes classifying different groups of obsessions, although the subtypes are not standardized or formally recognized. For example, “contamination OCD” groups together different obsessions of interacting with substances, such as a fear of HIV or asbestos (28), whereas “scrupulosity” groups together different obsessions of being amoral or sacrilegious. Some of these themes concern taboo topics, such as “sexual” or “pedophilic OCD.” These named subtypes can be helpful for individuals with OCD to connect with others with similar themes and for clinicians to have a better starting point of how to treat the individual (PhD, 2025).
@@ 3. Related Work
This mirrors the lack of research on OCD within HCI. Despite the lifetime prevalence of OCD at over 2%, OCD is rarely considered in the ACM literature. A search in the ACM digital library in July 2025 returned only nine publications with the term “Obsessive Compulsive Disorder” or “OCD” in the title, one of which uses the term “OCD” solely as a catchy pun. The remaining publications describe OCD (Wang, 2020), introduce technological approaches to OCD therapy (Wang et al., 2025; Schepers et al., 2016; Ma and Hu, 2025; Khan et al., 2011), work to understand deep brain stimulation for OCD (Eigen et al., 2004; Cohn et al., 2018), and propose virtual reality as a way to enhance empathy for those with OCD (Liaskou and Rizopoulos, 2023). Of those about OCD, only one is over five pages in length, and none explore the tensions between OCD and technology. This paper provides a necessary inspection of this tension within the realm of AI.
@@ 4. Method
In this paper, I analyzed 100 posts from a popular subreddit (sub) about OCD that allows research. Due to changes in the Reddit API, I used a manual search process for posts containing the keywords “AI” or “ChatGPT” on the sub. Posts were sorted by relevance. I removed duplicate posts from the dataset, as well as two posts stating within the text that the poster was under 18 and two posts that were not about AI. Finally, the next most relevant posts were included to reach 100 posts. The choice of 100 posts was sufficient to reach saturation, as evidenced by repeated themes and rich qualitative data.
I analyzed the data using a qualitative grounded theory process with constant comparison (Corbin and Strauss, 2026). I first analyzed the data using an inductive coding process, finding posts to be able to be categorized by whether they focus on describing a personal experience with AI-based obsessions, AI-based compulsions, or another topic. From there, AI-based obsession posts were categorized using a deductive process based on common OCD subtypes. AI-based compulsion posts were categorized inductively by compulsion type. Finally, the other posts were categorized inductively by type of conversation, such as warnings to others with OCD or moderator-led discussions. The data is presented in this paper with a focus on qualitative interpretation following the methodology in learning sciences by Hammer and Berland, rather than a quantification of codes and themes (Hammer and Berland, 2014).
@@ 5. Findings
The data can be split into three categories: posts where individuals share AI-related obsessions, AI-related compulsions, and ideas about AI and OCD more generally. Posts that fit into of these categories were spread relatively evenly in the dataset (minimum of 28 and maximum of 38 posts per category).
@@ 5.1. AI-Based Obsessions
Obsessions tend to cover a broad range of topics. I found this to be true in this dataset as well, despite all obsessions in this category being related to AI in some way. While some AI-related obsessions were repeated by multiple posters, such as a fear that AI will take over their job or creative pursuits (“I keep seeing posts of people being excited about AI art… but I’m terrified. I worry I might die of starvation because I am a musician– why would anyone pay me to make music when a bot can do it cheaper?”) and a fear of being inaccurately accused of submitting AI work (“I’m terrified of my English class b/c I hear about other students getting accused of cheating on their homework by using AI… I am just so anxious about this I might not sign up for school next semester.”), others were only seen once. A sample of five unique obsessions and their related subtype seen in the dataset are shown in Table 1. Notably, the obsessions are all able to be mapped to an existing subtype classification used by people in the OCD and mental health community. [^2] See: https://www.treatmyocd.com/education/different-types-of-ocd This is promising for clinical practice, as it suggests existing methods of treatment may be able to be applied to these novel AI-based obsessions.
@@
Scrupulosity (Moral and Religious) “i try so hard to not support corporations that i think are doing the wrong things. i try so hard to not use chatgpt, youtube, google, etc.– even switching to linux– but i get so stressed out about it. it’s really hard to avoid, and i feel so lonely, like i am the only one who cares about this stuff”
@@
Table 1. A sample of five distinct obsessions seen in the dataset.
@@ 5.2. AI-Based Compulsions
The AI-based compulsions seen in the dataset were more uniform than the obsessions. Despite having over 25 posts mentioning AI-based compulsions, posts can be classified into three types: reassurance-seeking, confession, and decision-making.
@@ 5.3. General Conversation
Many of the posts that fall outside the above two categories were posts warning others that they should not use GenAI for reassurance, such as one post titled, “STOP USING CHATGPT FOR REASSURANCE” (capitalization original). Other posts in this category considered using GenAI as a therapist to help structure Exposure and Response Preventon (ERP), the gold-standard behavioral treatment for OCD [^3] https://iocdf.org/ocd-treatment-guide/, or to help explain psychological concepts to themselves.
@@
Posts also mentioned a high frequency of use of GenAI, such as one that read, “I started by just asking ChatGPT for homework help, but then I started asking it questions about everything– what to wear, what to eat. Then I just go so used to using it that I started using it every single day.” One also noted the difference between using GenAI for reassurance-seeking compared to a basic web search, writing, “Googling will also give me reassurance, but it doesn’t get into the specifics the way ChatGPT does. I can ask it anything and get reassurance. It’s so hard to stop. It feels like an addiction.”
@@ 6.2. On Using GenAI for Heavy Disguise
6.3. Limitations and Future Work
This study does not claim to represent all possible interactions between OCD and AI, and future work should dive deeper into the subject, such as surveying a larger quantity of individuals with OCD to gather more variety in AI-based obsessions and compulsions, or interviewing clinicians who specialize in treating OCD to gain a deeper understanding of if and how AI has impacted the course of OCD or its treatment.
7. Conclusion
In this paper, I identified novel manifestations of OCD through AI-based obsessions and compulsions. I argue that GenAI’s function as a Reassurance Robot is harmful to individuals with OCD, and future GenAI design must take this into account by creating GenAI that does not provide these accommodations.

### v2

# arXiv:2602.19401v2[cs.HC] 03 Mar 2026

@@ Abstract.
Obsessive Compulsive Disorder (OCD) is a mental health disorder characterized by distressing repetitive patterns of thought, called obsessions, and behaviors aimed to reduce the distress, called compulsions. The explosion of artificial intelligence into the modern zeitgeist through the introduction of generative AI (GenAI) systems such as ChatGPT has led to novel obsessions and compulsions involving AI in individuals with OCD. Through an exploratory qualitative analysis of 100 Reddit posts related to AI on a popular subreddit for OCD, I examine ways AI is impacting the presentation of OCD, including novel examples of AI-based obsessions and compulsions. I argue that GenAI in its current form harms individuals with OCD by becoming “Reassurance Robots,” and that future designs of GenAI must take OCD into account. I recommend further work explore the intersection between OCD and GenAI.
@@
$c$s: Human-centered computing Accessibility theory, concepts and paradigmsccs: Social and professional topics People with disabilitiesccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Despite having a lifetime prevalence of 2.3% of the population in the United States, with approximately 50% of those with the disorder reporting serious impairment (17), there is a major gap in the diagnosis and treatment of individuals with OCD. The mean difference between age of symptom onset and age of diagnosis is often over ten years (Ziegler et al., 2021), and a recent report found that up to 75% of cases of OCD in the United States are undetected or undiagnosed (Deusser et al., 2025), leaving a significant population with OCD symptoms that are currently untreated. As such, it is imperative that the disorder receive more attention and research to improve outcomes in individuals living with OCD.
With the advent of GenAI into the popular zeitgeist, it has been reported that GenAI has become a topic of obsessions and a site of compulsions for people with OCD (i.e. (Samuel, 2025)). At a time when many researchers are looking to GenAI to help diagnose or treat OCD (Kim et al., 2025), there needs to be an understanding of how it can also cause harm. At this time, most research on the impact of GenAI on mental health has been in the context of depression and suicidality (Amine et al., 2025), however, given the high prevalence of OCD, it is critical that research study the impact of GenAI on OCD as well.
This paper contributes what I believe to be the first publication exploring the ways in which GenAI has led to novel presentations of OCD through AI-based obsessions and compulsions. In the discussion, I argue that the current design of GenAI harms individuals with OCD by becoming “Reassurance Robots,” that future designs of GenAI should design defensively against this, and that further research must be conducted on the subject.
@@ 2. Background
OCD is often thought of as broken down into subtypes classifying different groups of obsessions, although the subtypes are not standardized or formally recognized. For example, “contamination OCD” groups together different obsessions of interacting with substances, such as a fear of blood or asbestos (29), whereas “scrupulosity” groups together different obsessions of being amoral or sacrilegious. Some of these themes concern taboo topics, such as “sexual” or “pedophilic OCD.” These named subtypes can be helpful for individuals with OCD to connect with others with similar themes and for clinicians to have a better starting point of how to treat the individual (PhD, 2025).
@@
A diagram showing a cycle. Obsession points to feeling of distress points to compulsion points to temporary relief points to obsession again.
@@ 3. Related Work
Despite the lifetime prevalence of OCD at over 2%, OCD is rarely considered in the ACM literature. A search in the ACM digital library in July 2025 returned only nine publications with the term “Obsessive Compulsive Disorder” or “OCD” in the title, one of which uses the term “OCD” solely as a catchy pun. The remaining publications describe OCD (Wang, 2020), introduce technological approaches to OCD therapy (Wang et al., 2025; Schepers et al., 2016; Ma and Hu, 2025; Khan et al., 2011), work to understand deep brain stimulation for OCD (Eigen et al., 2004; Cohn et al., 2018), and propose virtual reality as a way to enhance empathy for those with OCD (Liaskou and Rizopoulos, 2023). Of those about OCD, only one is over five pages in length, and none explore the tensions between OCD and technology. This paper provides a necessary inspection of this tension within the realm of GenAI.
@@ 4. Method
In this paper, I analyzed 100 posts from a popular subreddit (sub) about OCD that allows research. Due to changes in the Reddit API, I used a manual search process for posts containing the keywords “AI” or “ChatGPT” on the sub. Posts were sorted by relevance. Skipping two posts stating within the text that the poster was under 18 and two posts that were not about AI, the first 100 unique posts were taken to form the dataset. The choice of 100 posts was sufficient to reach saturation, as evidenced by repeated themes and rich qualitative data. Data was collected in January 2025. Most posts (n=60) were from 2025, with an additional 1 from 2021, 3 from 2022, 14 from 2023, 15 from 2024, and 7 from 2026.
I analyzed the body of the posts using a qualitative grounded theory process with constant comparison (Corbin and Strauss, 2026). I first analyzed the data using an inductive coding process, finding posts to be able to be categorized by whether they focus on describing a personal experience with AI-based obsessions, AI-based compulsions, or another topic. From there, AI-based obsession posts were categorized using a deductive process based on common OCD subtypes. [^1] See: https://www.treatmyocd.com/education/different-types-of-ocd AI-based compulsion posts were categorized inductively by compulsion type (reassurance-seeking, confession, and decision making appeared in the dataset). Finally, the other posts were categorized inductively by type of conversation, such as warnings to others with OCD or moderator-led discussions. The data is presented in this paper with a focus on qualitative interpretation following the methodology in learning sciences by Hammer and Berland, rather than a quantification of codes and themes (Hammer and Berland, 2014). Given the use of a single author annotator and non-random sample, this data should be considered an exploration of the topic, with further research needed.
@@ 5. Findings
The data can be split into three categories: posts where individuals share AI-related obsessions, AI-related compulsions, and ideas about AI and OCD more generally. Posts that fit into of these categories were spread relatively evenly in the dataset, with 35 posts categorized as describing AI-based obsessions, 31 as AI-based compulsions, and 38 as other conversation (posts could be in multiple categories).
@@ 5.1. AI-Based Obsessions
Obsessions tend to cover a broad range of topics. I found this to be true in this dataset as well, even though all obsessions in this category being related to AI in some way. While some AI-related obsessions were repeated by multiple posters, such as an existential fear that AI will take over their job or creative pursuits (“I keep seeing posts of people being excited about AI art… but I’m terrified. I worry I might die of starvation because I am a musician– why would anyone pay me to make music when a bot can do it cheaper?”) and a fear of being inaccurately accused of submitting AI work (“I’m terrified of my English class b/c I hear about other students getting accused of cheating on their homework by using AI… I am just so anxious about this I might not sign up for school next semester.”), others were only seen once. Many posts could be mapped to existing known subtypes as shown in Table 1, although others may represent new types of OCD such as the fear of being accused of using AI as stated above. The most common obsession could be categorized as existential OCD, with almost half (n=16, 46%) of all AI-based obsessions relating to this subtype, such as the poster who wrote, “I’ve been reading the news, and I’m worried AI is going to destroy the world.”
@@
Scrupulosity (Moral or Religious) “i try so hard to not support corporations that i think are doing the wrong things. i try so hard to not use chatgpt, youtube, google, etc.– even switching to linux– but i get so stressed out about it. it’s really hard to avoid, and i feel so lonely, like i am the only one who cares about this stuff”
@@
Table 1. A sample of five distinct obsessions from the dataset.
@@ 5.2. AI-Based Compulsions
The AI-based compulsions seen in the dataset were more uniform than the obsessions. With over 25 posts mentioning AI-based compulsions, posts can be classified into three types: reassurance-seeking, confession, and decision-making.
@@ 5.3. General Conversation
Many of the posts that fall outside the above two categories were posts warning others that they should not use GenAI for reassurance, such as one post titled, “STOP USING CHATGPT FOR REASSURANCE” (capitalization original). Other posts in this category considered using GenAI as a therapist to help structure Exposure and Response Preventon (ERP), the gold-standard behavioral treatment for OCD (18), or to help explain psychological concepts to themselves.
@@
Posts also mentioned a high frequency of use of GenAI, such as one that read, “I started by just asking ChatGPT for homework help, but then I started asking it questions about everything– what to wear, what to eat. Then I just got so used to using it that I started using it every single day.” One also noted the difference between using GenAI for reassurance-seeking compared to a basic web search, writing, “Googling will also give me reassurance, but it doesn’t get into the specifics the way ChatGPT does. I can ask it anything and get reassurance. It’s so hard to stop. It feels like an addiction.”
@@ 6.2. On Using GenAI for Heavy Disguise
6.3. Limitations, Future Work, and Conclusion
Despite the high prevalence of OCD and the significant impairment that it often causes (17), there is very little research conducted within computing on OCD. This exploratory paper shows that technology such as GenAI can have an impact on the mental health of those with OCD. Due to limitations in this paper such as its reliance on self-reported data and diagnosis, the use of a single data annotator for the qualitative analysis, and using the top posts as sorted by relevance, this study should be considered an exploration into the topic to promote further research and discussion. Further research should be conducted to create a deeper understanding of the impacts of technology on OCD, such as surveying a larger quantity of individuals with OCD to gather more variety in AI-based obsessions and compulsions, or interviewing clinicians who specialize in treating OCD to gain an understanding of if and how AI has impacted the course of OCD or its treatment.
In this paper, I identified novel manifestations of OCD through AI-based obsessions and compulsions. I argue that GenAI’s function as a Reassurance Robot is harmful to individuals with OCD, and future GenAI design must take this into account by creating GenAI that does not provide these accommodations. Future work within computing should continue to explore technology’s impact on OCD.
Acknowledgements.
Thank you to the anonymous reviewers who provided important feedback on the original manuscript. No AI was used in the creation of this paper.
I am deeply grateful to those who have shared their OCD stories with me and those who have heard mine.